OpenStudy (anonymous):

A die is rolled. What is the probability of rolling the following? a multiple of 3 or a multiple of 4

OpenStudy (anonymous):

only multiples of 3 on one die are 3 and 6

OpenStudy (anonymous):

A die ir rolled therefore sample space S={1,2,3,4,5,6} i.e. n(s) =6 Let A be the event of getting a multiple of 3 and B be the event of getting a multiple of 4 A={3,6} i.e. n(A)=2 & B={4} i.e. n(B)=1 therefore Probability of getting a multiple of 3 \[P(A)= \frac{n(A)}{n(S)} = \frac{2}{6} = \frac{1}{3} \] Probability of getting a multiple of 4 \[P(B)= \frac{n(B)}{n(S)} = \frac{1}{6} \]